# Deep Research Agents: A Systematic Examination And Roadmap

### v1

@@ 3.3.4 Memory Mechanism for Long-Context Optimization
Memory mechanisms empower DR agents to persistently capture, organize, and recall relevant information across multiple retrieval rounds, thereby reducing redundant queries and improving both the efficiency and coherence of DR tasks. During the DR process, agents typically perform extensive multi-round retrieval, generating hundreds of thousands of tokens (or even millions). Although recent advances in LLMs have significantly expanded context window sizes, current limits still constrain tasks involving extremely long contexts. To address these challenges, DR systems have implemented various optimizations for processing extended contexts. Broadly, these optimizations can be categorized into three main strategies: (i) Expanding the Context Window Length; (ii) Compressing Intermediate Steps; (iii) Utilizing External Structured Storage for Temporary Results.
@@
Extending the context window length is the most intuitively effective approach, exemplified by Google’s Gemini model geminideepresearch , which supports a context window of up to one million tokens, supplemented by a RAG setup. Despite its straightforwardness, this method often incurs high computational costs and may lead to inefficiencies in resource utilization during practical deployments.
@@
An alternative strategy involves compressing or summarizing intermediate reasoning steps, significantly reducing the number of tokens processed by the model and thereby improving both efficiency and output quality. Representative frameworks such as The AI Scientist lu2024aiscientistfullyautomated and CycleResearcher weng2024cycleresearcher pass summarized intermediate results between workflow phases. Further, Li et al. li2025search introduced the concept of “Reason-in-Documents,” utilizing LRMs to compress documents, substantially reducing token volume and enhancing model decision-making efficiency. However, a potential drawback of this approach is the loss of detailed information, potentially impacting the precision of subsequent reasoning.
@@
Utilizing external structured storages for preserving and retrieving historical information enables DR agents to persistently and efficiently store vast amounts of past context beyond the constraints of the context window, improving memory capacity, retrieval speed, and semantic relevance. Popular open-source frameworks such as Manus manus2025 , OWL owl2025 , Open Manus openmanus2025 , and Avatar wu2024avataroptimizingllmagents utilize external file systems to store intermediate outcomes and historical data for subsequent retrieval. Frameworks like WebThinker Li2025webthinker and AutoAgent tang2025autoagentfullyautomatedzerocodeframework have developed self-managing modules that leverage vector databases to support scalable memory storage and fast similarity-based lookup. Beyond plain text or vector stores, some works propose more semantically structured memory frameworks: for instance, Wu et al. wu2025agentic employ knowledge graphs to capture intermediate reasoning processes and thereby enhance the precision of information reuse, while Agentrxiv schmidgall2025agentrxiv simulates an academic repository akin to arXiv for storing and retrieving relevant outcomes from other agents. Although these structured approaches offer superior semantic retrieval efficiency and accuracy, they typically entail higher development and maintenance costs due to the need for meticulous data structure design and management.
@@ 3.4 Tuning: Beyond Prompting toward Capability Enhancement
ReSearch chen2025learning GRPO Qwen2.5-7B-Inst, Qwen2.5-32B-Inst 2WikiMultiHopQA Rule-Outcome Mar-2025
R1-Searcher song2025r1 Reinforce++, GRPO Qwen2.5-7B-InSt, LLaMA-3.1-8B-Inst 2WikiMultiHopQA, HotpotQA Rule-Outcome Mar-2025
Search-R1 jin2025search PPO, GRPO Qwen2.5-3B/7B, LLaMA3.2-3B-Inst NQ, HotpotQA Rule-Outcome Mar-2025
DeepResearcher zheng2025deepresearcherscalingdeepresearch GRPO Qwen2.5-7B-Inst NQ, HotpotQA Rule-Outcome Apr-2025
@@
WebThinker Li2025webthinker Iterative Online DPO QwQ-32B Expert Dataset Rule-Outcome Apr-2025
@@
SimpleDeepSearcher SimpleDeepSearcher PPO Qwen-2.5-7B-In, Qwen-2.5-32B-In, Deepseek-Distilled-Qwen-32B, QwQ-32B NQ, HotpotQA, 2WikiMultiHopQA, Musique, SimpleQA, MultiHop-RAG Process-based reward Apr-2025
PANGU DEEPDIVER shi2025pangu GRPO Pangu-7B-Reasoner WebPuzzle Rule-Outcome May-2025
@@
Prompt-based methods directly leverage the capabilities of pre-trained LLMs, enabling complex functionalities without expensive fine-tuning or additional training. However, it remains challenging to systematically optimize prompt structures and workflows. Moreover, since an agent’s performance is inherently limited by its backbone LLM, increasing the complexity of decision-making processes quickly reaches the model’s performance ceiling. To overcome these limitations, it is essential to incorporate advanced optimization techniques such as fine-tuning, reinforcement learning (RL) or hybrid training paradigms to further extend the model’s inherent capabilities. Below, we discuss the two main tuning paradigms, supervised fine-tuning (SFT) and RL, and highlight how each extends agent capabilities beyond prompt-only methods.
@@ 3.4.1 SFT-based Optimization
Prompt-based approaches, while effective for rapid adaptation, are fundamentally constrained by the intrinsic generalization capacity of backbone LLMs and often exhibit limited robustness in complex task settings. In order to address these limitations, researchers have increasingly explored fine-tuning methodologies aimed at systematically optimizing LLMs for critical components of deep research agents. These components include search query formulation, structured report generation, and external tool utilization. These efforts aim to enhance retrieval quality, mitigate hallucinations, and enable more reliable long-form and evidence-grounded generation.
@@
An early milestone in this research direction is Open-RAG islam2024open , which augments data construction with diverse supervisory signals, including retrieval tokens, relevance tokens, grounding tokens, and utility tokens. Through adversarial training, Open-RAG improves the model’s capability to filter irrelevant information, thereby enhancing both retrieval accuracy and the quality of downstream tasks. Building upon this foundation, AUTO-RAG yu2024auto enhances the autonomous iterative retrieval capabilities of LLMs. In contrast to earlier multi-hop retrieval approaches that relied on few-shot prompting or hand-crafted templates jiang2023active ; feng2024retrieval ; wang2024llms , AUTO-RAG constructs reasoning-grounded instruction datasets, enabling models to autonomously plan retrieval queries and engage in multi-round interactions with retrievers. The model dynamically refines its retrieval strategy during generation, gathering sufficient evidence before synthesizing a final answer. Extending these retrieval-centric innovations, DeepRAG guan2025deeprag proposes a binary tree search mechanism that recursively generates sub-queries and constructs multi-turn retrieval trajectories. This mechanism enables the model to judiciously balance between internal parametric knowledge and external retrieval-based rollouts. Consequently, it enhances search efficiency and mitigates redundant external queries.
@@
In order to further reduce reliance on manually constructed supervised fine-tuning (SFT) datasets, recent work has sought to reduce dependence on manually constructed supervised fine-tuning datasets by developing fine-tuning strategies based on rejection sampling. CoRAG wang2025chainofretrievalaugmentedgeneration uses rejection sampling to extract intermediate retrieval chains from standard question answering datasets, allowing for stepwise retrieval augmentation and dynamic reformulation of subqueries as context evolves instead of supervising only final outputs. Li et al. li2025start propose a hint-infer mechanism that monitors token patterns during generation and triggers external computational tools, such as Python executors or hint libraries when specific cues are detected. After an initial supervised fine-tuning phase, the model undergoes a rejection sampling fine-tuning process that teaches it to generate its own prompts and invoke tools autonomously without reliance on hand-curated demonstrations. ATLAS chen2025atlas proposes a novel approach for LLM-based agents that trains exclusively on selected critical steps from expert trajectories, significantly improving generalization performance.
@@
Although these SFT methods enhance the generalization of deep research agents by supporting dynamic retrieval planning, structured information synthesis, and integrated tool use, they remain confined to offline, static retrieval pipelines characteristic of retrieval-augmented systems. In contrast, reinforcement learning offers a more adaptive solution for online query generation and tool invocation. By learning from real-time reward signals, reinforcement learning agents acquire the ability to formulate effective search queries and determine the optimal timing for tool calls. This approach addresses the limitations of synthetic demonstration data and distributional shifts, yielding more robust and adaptive performance in open-ended research environments.
3.4.2 Reinforcement Learning-based Optimization
@@
RL-based methods optimize DR agents by directly enhancing their adaptive capabilities and generalization across diverse tasks, surpassing conventional instruction-following or pattern learning approaches. Recent advances have demonstrated that end-to-end RL training significantly strengthens iterative information retrieval, dynamic tool invocation, and integrated reasoning capabilities within DR agents. See comparative analysis in Table 3.
@@
Early RL-based approaches such as DeepRetrieval jiang2025deepretrieval optimized query generation for improved information retrieval quality, effectively enhancing downstream text generation by producing more relevant search results. Building on query optimization, ReSearch chen2025learning extended RL to adaptive reasoning over retrieved information. The model dynamically refined search strategies and iteratively updated results based on continuous feedback, significantly improving task-solving accuracy. Subsequently, R1-Searcher song2025r1 further optimized retrieval interactions, explicitly training models to refine search strategies through carefully designed reward functions. This allowed better exploitation of external information and improved search result relevance.
@@
Search-R1 jin2025search advanced RL-based retrieval by structurally integrating sophisticated search interactions with complex reasoning processes. The method systematically bridged query generation and information reasoning, enabling nuanced responses through refined integration of retrieved content. Finally, this research line culminated in the development of Agent-R1 Agent-R1 , a comprehensive DR framework integrating RL into end-to-end training of LLM agents. Agent-R1 leveraged diverse tools such as APIs, search engines, and databases, achieving autonomous multi-step task execution and dynamic tool coordination. Through RL-driven optimization across its entire pipeline, Agent-R1 demonstrated advanced capabilities in adaptive planning, iterative execution, and task refinement. Moreover, WebThinker Li2025webthinker integrates a Web Explorer module for dynamic multi-hop web exploration and employs Iterative Online Direct Pre ference (DPO) Optimization to seamlessly interleave search, navigation, and report drafting during reasoning, while Pangu DeepDiver shi2025pangu builds on the 7B Pangu model pretrained on Huawei’s Ascend NPUs by introducing Search Intensity Scaling (SIS) through a two-phase SFT and RL curriculum, enabling adaptive adjustment of search depth and frequency in open-web environments.
@@
Table 3 reveals three key RL implementation patterns in DR systems: 1) Industrial systems like Gemini DR geminideepresearch and Grok DeepSearch grokdeepresearch employ proprietary RL implementations with undisclosed details, 2) Academic approaches chen2025learning ; song2025r1 favor modular RL optimization using GRPO shao2024deepseekmath and Reinforce++ hu2025reinforce with transparent reward designs, and 3) Emerging hybrid systems like SimpleDeepSearcher SimpleDeepSearcher combine process-based rewards with multi-task training across 6 QA datasets. The table also highlights the prevalence of Qwen2.5 and LLaMA3 model families as preferred base architectures for RL optimization.
@@
Reward Model and Policy Model. Most current open-source RL implementations of DR agents, including the methods discussed above, commonly adopt rule-based reward models that explicitly define task-specific objectives such as retrieval relevance, information accuracy, or successful tool invocation. To efficiently perform policy optimization, recent systems have increasingly utilized Proximal Policy Optimization (PPO) schulman2017proximal and Group Relative Policy Optimization (GRPO) shao2024deepseekmath . In particular, GRPO fundamentally reconfigures the advantage estimation paradigm by replacing traditional value functions with group-relative advantage computation. It expands reward space through intra-group normalization, sparse binary rewards are transformed into continuous advantage values spanning wider ranges. This expanded signal space provides richer gradient information for policy updates, as evidenced higher high-reward response density compared to PPO. In addition, GRPO provides a variance suppression mechanism by constraining advantage estimation within dynamically clustered response groups, such as grouping by reasoning depth or tool usage patterns, reducing policy gradient variance through local standardization. In contrast to PPO, GRPO eliminates separate value networks, removing conflicting optimization objectives between policy and value functions. Empirical measurements show GRPO reduces gradient direction conflicts from 12 to 3 per training epoch, significantly accelerating convergence. As a result, GRPO outperforms conventional PPO in wider reward distribution coverage, enhancing exploration capacity and faster KL divergence stabilization during alignment.

### v2

@@ 3.3.4 Memory Mechanism for Long-Context Optimisation
Memory mechanisms empower DR agents to persistently capture, organise, and recall relevant information across multiple retrieval rounds, thereby reducing redundant queries and improving both the efficiency and coherence of DR tasks. During the DR process, agents typically perform extensive multi-round retrieval, generating hundreds of thousands of tokens (or even millions). Although recent advances in LLMs have significantly expanded context window sizes, current limits still constrain tasks involving extremely long contexts. To address these challenges, DR systems have implemented various optimisations for processing extended contexts. Broadly, these optimisations can be categorised into three main strategies: (i) Expanding the Context Window Length; (ii) Compressing Intermediate Steps; (iii) Utilising External Structured Storage for Temporary Results.
@@
Extending the Context Window Length. It is the simple but intuitively effective approach, exemplified by Google’s Gemini model (geminideepresearch, ), which supports a context window of up to one million tokens, supplemented by a RAG setup. Despite its straightforwardness, this method often incurs high computational costs and may lead to inefficiencies in resource utilisation during practical deployments.
@@
Compressing Intermediate Step. An alternative strategy involves compressing or summarising intermediate reasoning steps, significantly reducing the number of tokens processed by the model and thereby improving both efficiency and output quality. Representative frameworks such as The AI Scientist (lu2024aiscientistfullyautomated, ) and CycleResearcher (weng2024cycleresearcher, ) pass summarised intermediate results between workflow phases. Further, Search-o1 (li2025search, ) introduced the concept of “Reason-in-Documents”, utilising LRMs to compress documents, substantially reducing token volume and enhancing model decision-making efficiency. Meanwhile, WebThinker (Li2025webthinker, ) uses an auxiliary model to compress the external information. However, a potential drawback of this approach is the loss of detailed information, potentially impacting the precision of subsequent reasoning.
@@
Utilising External Structured Storage. This is for preserving and retrieving historical information, enabling DR agents to persistently and efficiently store vast amounts of past context beyond the constraints of the context window, improving memory capacity, retrieval speed, and semantic relevance. Popular open-source frameworks such as Manus (manus2025, ), OWL (owl2025, ), Open Manus (openmanus2025, ), and Avatar (wu2024avataroptimizingllmagents, ) utilise external file systems to store intermediate outcomes and historical data for subsequent retrieval. Frameworks like AutoAgent (tang2025autoagentfullyautomatedzerocodeframework, ) have developed self-managing modules that leverage vector databases to support scalable memory storage and fast similarity-based lookup. Beyond plain text or vector stores, some works propose more semantically structured memory frameworks: for instance, Agentic Reasoning (wu2025agentic, ) employ knowledge graphs to capture intermediate reasoning processes and thereby enhance the precision of information reuse, while Agentrxiv (schmidgall2025agentrxiv, ) simulates an academic repository akin to arXiv for storing and retrieving relevant outcomes from other agents. Furthermore, Agent-KB (tang2025agent, ) and Alita (qiu2025alita, ) construct shared knowledge bases and optimised toolsets for agentic problem-solving. Although these structured approaches offer superior semantic retrieval efficiency and accuracy, they typically entail higher development and maintenance costs due to the need for meticulous data structure design and management.
@@ 3.4 Tuning: Beyond Prompting toward Capability Enhancement
Agentic Reasoning wu2025agentic DeepSeek-R1, Qwen2.5 GPQA Rule-Outcome Feb-2025
Towards an AI co-scientist Gemini 2.0 Feb-2025
@@
ReSearch chen2025learning GRPO shao2024deepseekmath Qwen2.5-7B-Inst, Qwen2.5-32B-Inst 2WikiMultiHopQA Rule-Outcome Mar-2025
R1-Searcher song2025r1 Reinforce++ hu2025reinforce , GRPO shao2024deepseekmath Qwen2.5-7B-InSt, LLaMA-3.1-8B-Inst 2WikiMultiHopQA, HotpotQA Rule-Outcome Mar-2025
Search-R1 jin2025search PPO schulman2017proximal , GRPO shao2024deepseekmath Qwen2.5-3B/7B, LLaMA3.2-3B-Inst NQ, HotpotQA Rule-Outcome Mar-2025
Nouswise (nouswise2025, ) Nouswise Mar-2025
DeepResearcher zheng2025deepresearcherscalingdeepresearch GRPO shao2024deepseekmath Qwen2.5-7B-Inst NQ, HotpotQA Rule-Outcome Apr-2025
@@
WebThinker Li2025webthinker Iterative Online DPO (rafailov2023direct, ) QwQ-32B Expert Dataset Rule-Outcome Apr-2025
@@
SimpleDeepSearcher SimpleDeepSearcher PPO schulman2017proximal Qwen-2.5-7B-In, Qwen-2.5-32B-In, Deepseek-Distilled-Qwen-32B, QwQ-32B NQ, HotpotQA, 2WikiMultiHopQA, Musique, SimpleQA, MultiHop-RAG Process-based reward Apr-2025
PANGU DEEPDIVER shi2025pangu GRPO shao2024deepseekmath Pangu-7B-Reasoner WebPuzzle Rule-Outcome May-2025
Tool-Star (dong2025tool, ) GRPO shao2024deepseekmath Qwen-2.5 NuminaMath, HotpotQA, 2WikiMultiHopQA Rule-Outcome May-2025
WebDancer (wu2025webdancer, ) DAPO (yu2025dapo, ) Qwen-2.5-7B/32B, QwQ-32B, DeepSeek-R1, GPT-4o CRAWLQA, E2HQA Rule-Outcome May-2025
O-agents (zhu2025oagents, ) GPT-4o, GPT-4.1, Claude-3.7-Sonnet, DeepSeek-R1, Gemini-2.5 Jun-2025
Kimi-Researcher (kimi2025researcher, ) REINFORCE Kimi k1.5/k2 Rule-Outcome Jun-2025
WebSailor (li2025websailor, ) DUPO Qwen-2.5-3B/7B/32B/72B SailorFog-QA Rule-Outcome Jul-2025
Agent-KB (tang2025agent, ) GPT-4o, GPT-4.1, Claude-3.7-Sonnet, o3-mini, Qwen-3 32B, DeepSeek-R1 Jul-2025
WebShaper (tao2025webshaper, ) GRPO shao2024deepseekmath Qwen-2.5-3B/7B/32B/72B, QwQ-32B WebShaper Rule-Outcome Jul-2025
Cognitive Kernel-Pro (wan2025cognitive, ) Claude-3.7-Sonnet, CK-Pro-8B OpenWebVoyager, Multi-hop URLQA, AgentWebQA, WebWalkerQA, DocBench, TableBench Aug-2025
WebWatcher (geng2025webwatcher, ) GRPO shao2024deepseekmath Qwen-2.5-VL-32B BrowseComp-VL, Long-tail VQA, Hard VQA Rule-Outcome Aug-2025
MiroRL (2025mirorl, ) GRPO shao2024deepseekmath Qwen3-14B MiroRL-GenQA Rule-Outcome Aug-2025
@@
Prompt-based methods directly leverage the capabilities of pre-trained LLMs, enabling complex functionalities without expensive fine-tuning or additional training. However, it remains challenging to systematically optimise prompt structures and workflows. Moreover, since an agent’s performance is inherently limited by its backbone LLM, increasing the complexity of decision-making processes quickly reaches the model’s performance ceiling. To overcome these limitations, it is essential to incorporate advanced optimisation techniques such as fine-tuning, reinforcement learning (RL) or hybrid training paradigms to further extend the model’s inherent capabilities. Below, we discuss the two main tuning paradigms, supervised fine-tuning (SFT) and RL, and highlight how each extends agent capabilities beyond prompt-only methods.
@@ 3.4.1 SFT-based Optimization
Prompt-based approaches, while effective for rapid adaptation, are fundamentally constrained by the intrinsic generalisation capacity of backbone LLMs and often exhibit limited robustness in complex task settings. In order to address these limitations, researchers have increasingly explored fine-tuning methodologies aimed at systematically optimising LLMs for critical components of deep research agents. These components include search query formulation, structured report generation, and external tool utilisation. These efforts aim to enhance retrieval quality, mitigate hallucinations, and enable more reliable long-form and evidence-grounded generation.
@@
An early milestone in this research direction is Open-RAG (islam2024open, ), which augments data construction with diverse supervisory signals, including retrieval tokens, relevance tokens, grounding tokens, and utility tokens. Through adversarial training, Open-RAG improves the model’s capability to filter irrelevant information, thereby enhancing both retrieval accuracy and the quality of downstream tasks. Building upon this foundation, AUTO-RAG (yu2024auto, ) enhances the autonomous iterative retrieval capabilities of LLMs. In contrast to earlier multi-hop retrieval approaches that relied on few-shot prompting or hand-crafted templates (jiang2023active, ; feng2024retrieval, ; wang2024llms, ), AUTO-RAG constructs reasoning-grounded instruction datasets, enabling models to autonomously plan retrieval queries and engage in multi-round interactions with retrievers. The model dynamically refines its retrieval strategy during generation, gathering sufficient evidence before synthesising a final answer. Extending these retrieval-centric innovations, DeepRAG (guan2025deeprag, ) proposes a binary tree search mechanism that recursively generates sub-queries and constructs multi-turn retrieval trajectories. This mechanism enables the model to judiciously balance between internal parametric knowledge and external retrieval-based rollouts. Consequently, it enhances search efficiency and mitigates redundant external queries.
@@
In order to further reduce reliance on manually constructed supervised fine-tuning (SFT) datasets, recent work has sought to reduce dependence on manually constructed supervised fine-tuning datasets by developing fine-tuning strategies based on rejection sampling. CoRAG (wang2025chainofretrievalaugmentedgeneration, ) uses rejection sampling to extract intermediate retrieval chains from standard question answering datasets, allowing for stepwise retrieval augmentation and dynamic reformulation of subqueries as context evolves instead of supervising only final outputs. Li et al. (li2025start, ) propose a hint-infer mechanism that monitors token patterns during generation and triggers external computational tools, such as Python executors or hint libraries, when specific cues are detected. After an initial supervised fine-tuning phase, the model undergoes a rejection sampling fine-tuning process that teaches it to generate its own prompts and invoke tools autonomously without reliance on hand-curated demonstrations. ATLAS (chen2025atlas, ) proposes a novel approach for LLM-based agents that trains exclusively on selected critical steps from expert trajectories, significantly improving generalisation performance.
@@
Although these SFT methods enhance the generalisation of deep research agents by supporting dynamic retrieval planning, structured information synthesis, and integrated tool use, they remain confined to offline, static retrieval pipelines characteristic of retrieval-augmented systems. In contrast, reinforcement learning offers a more adaptive solution for online query generation and tool invocation. By learning from real-time reward signals, reinforcement learning agents acquire the ability to formulate effective search queries and determine the optimal timing for tool calls. This approach addresses the limitations of synthetic demonstration data and distributional shifts, yielding more robust and adaptive performance in open-ended research environments.
3.4.2 Reinforcement Learning-based Optimisation
@@
RL-based methods optimise DR agents by directly enhancing their adaptive capabilities and generalisation across diverse tasks, surpassing conventional instruction-following or pattern learning approaches. Recent advances have demonstrated that end-to-end RL training significantly strengthens iterative information retrieval, dynamic tool invocation, and integrated reasoning capabilities within DR agents. See comparative analysis in Table LABEL:tab:tuning.
@@
Early RL-based approaches such as DeepRetrieval (jiang2025deepretrieval, ) optimised query generation for improved information retrieval quality, effectively enhancing downstream text generation by producing more relevant search results. Building on query optimisation, ReSearch (chen2025learning, ) extended RL to adaptive reasoning over retrieved information. The model dynamically refined search strategies and iteratively updated results based on continuous feedback, significantly improving task-solving accuracy. Subsequently, R1-Searcher (song2025r1, ) further optimised retrieval interactions, explicitly training models to refine search strategies through carefully designed reward functions. This allowed better exploitation of external information and improved search result relevance.
@@
Search-R1 (jin2025search, ) advanced RL-based retrieval by structurally integrating sophisticated search interactions with complex reasoning processes. The method systematically bridged query generation and information reasoning, enabling nuanced responses through refined integration of retrieved content. Finally, this research line culminated in the development of Agent-R1 (Agent-R1, ), a comprehensive DR framework integrating RL into end-to-end training of LLM agents. Agent-R1 leveraged diverse tools such as APIs, search engines, and databases, achieving autonomous multi-step task execution and dynamic tool coordination. Through RL-driven optimisation across its entire pipeline, Agent-R1 demonstrated advanced capabilities in adaptive planning, iterative execution, and task refinement. Moreover, WebThinker (Li2025webthinker, ) integrates a Web Explorer module for dynamic multi-hop web exploration and employs Iterative Online Direct Preference Optimisation (DPO) to seamlessly interleave search, navigation, and report drafting during reasoning, while Pangu DeepDiver (shi2025pangu, ) builds on the 7B Pangu model pretrained on Huawei’s Ascend NPUs by introducing Search Intensity Scaling (SIS) through a two-phase SFT and RL curriculum, enabling adaptive adjustment of search depth and frequency in open-web environments.
@@
Table LABEL:tab:tuning reveals three key RL implementation patterns in DR systems: 1) Industrial systems like Gemini DR (geminideepresearch, ) and Grok DeepSearch (grokdeepresearch, ) employ proprietary RL implementations with undisclosed details, 2) Academic approaches (chen2025learning, ; song2025r1, ) favor modular RL optimization using GRPO (shao2024deepseekmath, ) and Reinforce++ (hu2025reinforce, ) with transparent reward designs, and 3) Emerging hybrid systems like SimpleDeepSearcher (SimpleDeepSearcher, ) combine process-based rewards with multi-task training across 6 QA datasets. The table also highlights the prevalence of Qwen2.5 and LLaMA3 model families as preferred base architectures for RL optimisation.
@@
Reward Model and Policy Model. Most current open-source RL implementations of DR agents, including the methods discussed above, commonly adopt rule-based reward models that explicitly define task-specific objectives such as retrieval relevance, information accuracy, or successful tool invocation. To efficiently perform policy optimisation, recent systems have increasingly utilised Proximal Policy Optimisation (PPO) (schulman2017proximal, ) and Group Relative Policy Optimisation (GRPO) (shao2024deepseekmath, ). In particular, GRPO fundamentally reconfigures the advantage estimation paradigm by replacing traditional value functions with group-relative advantage computation. It expands reward space through intra-group normalisation, and sparse binary rewards are transformed into continuous advantage values spanning wider ranges. This expanded signal space provides richer gradient information for policy updates, as evidenced higher high-reward response density compared to PPO. In addition, GRPO provides a variance suppression mechanism by constraining advantage estimation within dynamically clustered response groups, such as grouping by reasoning depth or tool usage patterns, reducing policy gradient variance through local standardisation. In contrast to PPO, GRPO eliminates separate value networks, removing conflicting optimisation objectives between policy and value functions. Empirical measurements show GRPO reduces gradient direction conflicts from 12 to 3 per training epoch, significantly accelerating convergence. As a result, GRPO outperforms conventional PPO in wider reward distribution coverage, enhancing exploration capacity and faster KL divergence stabilisation during alignment.